OpenStudy (anonymous):

Bella Thorne here just asking a question, if it takes 5 men to build a wall in 3 days how much does it take to build in 10 days

OpenStudy (jdoe0001):

w0t?

undeadknight26 (undeadknight26):

lol imagine if this was the real bella thorne... any who...this should be easy....

undeadknight26 (undeadknight26):

yet to lazy to answer it...

OpenStudy (wolf1728):

5 men build a wall in 3 days 1 man can build a wall in 15 days 2 men can build a wall in 7.5 days Do you want a fractional answer?

OpenStudy (wolf1728):

In one day, one man can build 1/15 of a wall. In 10 days, one man can build 2/3 of a wall. In 10 days, 1 and 1/3 men can build one wall.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

lol bella thorne. I met the real one :D

OpenStudy (anonymous):

rate of one man is 1/(5*3) = 1/15 so you have m (1/15) (10) = 1 m = 1.5 it takes 1 and a half men

OpenStudy (jdoe0001):

" 1 and a half men" <--- so one man and a midget =)

OpenStudy (jdoe0001):

unless is 1 man and half.... but what half... the wait up or down? I'd think the waist up, since that one has arms, he'd need a skateboard to move around though

OpenStudy (anonymous):

one men and a midget works fine lol